環境システム工学特論 2 単位 (選択)

Advanced Environmental Systems Engineering  木戸口善行 · 教授/ 環境創生工学専攻エコシステム工学コース資源循環工学講座

上月康則 · 教授/ 環境創生工学専攻エコシステム工学コース社会環境システム工学講座，近藤光男 · 教授/ 環境創生工学専攻エコシステム工学コース社会環境システム工学講座橋本修一 · 教授/ 環境創生工学専攻エコシステム工学コース資源循環工学講座，藤澤正一郎 · 教授/ 環境創生工学専攻エコシステム工学コース社会環境システム工学講座奥嶋政嗣 · 准教授/ 環境創生工学専攻エコシステム工学コース社会環境システム工学講座，松尾繁樹 · 准教授/ 環境創生工学専攻エコシステム工学コース資源循環工学講座山中亮一 · 講師/ 環境創生工学専攻エコシステム工学コース社会環境システム工学講座，富田卓朗 · 助教/ 環境創生工学専攻エコシステム工学コース資源循環工学講座佐藤克也 · 講師/ 環境創生工学専攻エコシステム工学コース社会環境システム工学講座，伊藤伸一 · 助教/ 環境創生工学専攻エコシステム工学コース社会環境システム工学講座

【授業目的】最近の市民生活環境一般に関する現状，問題点を理解し，それらの 問題点に関してエコロジーの立場から解決すべく行われている最新の研究と 産業界での取り組みを理解する． 【授業概要】市民生活と環境の問題について理解し，その工学的視野を広げるこ とを目的とする．環境システムの構成要因と構造を理解させ，市民生活や防 災とシステムとの関係，さらに技術革新や政策によるシステムの変化につい ての事例を交えて講述する． 【授業形式】講義 【キーワード】環境政策，自然災害，省エネ技術，福祉工学，マイクロ機械工学 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【履修要件】 [要件] 【履修上の注意】 [注意] 【到達目標】 1. 環境システムの構成要因の理解 2. 環境システムの構造の理解 3. 市民生活と防災システムの関係の理解 【授業計画】 1. 政策と防災システムの関係 1 2. 政策と防災システムの関係 2 3. 環境システムの構成 4. 環境システムの構成要因の事例紹介 5. 環境システムの構造 6. 環境システムの構造の事例紹介 7. 市民生活とナノテクノロジー 1 8. 市民生活とナノテクノロジー 2 9. レポート中間発表 10. 福祉工学 1

11. 福祉工学 2

12. 生態系工学 1

13. 生態系工学 2

14. 水素エンジン

15. ディーゼルエンジン【成績評価基準】授業時間中に課すレポートで評価する【教科書】授業中に紹介する【参考書】授業中に紹介する【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216598 【対象学生】 [対象学生] 【連絡先】

⇒ 佐藤 (エコ棟 705, 088-656-2168, katsuyas@eco.tokushima-u.ac.jp) ����

Advanced Environmental Systems Engineering 2 units (selection)

Yoshiyuki Kidoguchi · Professor / Resource Circulatory Engineering, Ecosystem Engineering, Earth and Life Environmental Engineering

Yasunori Kozuki · Professor / Social Environment Systems Engineering, Ecosystem Engineering, Earth and Life Environmental Engineering

, Akio Kondo · Professor / Social Environment Systems Engineering, Ecosystem Engineering, Earth and Life Environmental Engineering Shuichi Hashimoto · Professor / Resource Circulatory Engineering, Ecosystem Engineering, Earth and Life Environmental Engineering

Shoichiro Fujisawa · Professor / Social Environment Systems Engineering, Ecosystem Engineering, Earth and Life Environmental Engineering Masashi Okushima · Associate Professor / Social Environment Systems Engineering, Ecosystem Engineering, Earth and Life Environmental Engineering

Shigeki Matsuo · Associate Professor / Resource Circulatory Engineering, Ecosystem Engineering, Earth and Life Environmental Engineering Ryoichi Yamanaka · Associate Professor / Social Environment Systems Engineering, Ecosystem Engineering, Earth and Life Environmental Engineering

Takuro Tomita · Assistant Professor / Resource Circulatory Engineering, Ecosystem Engineering, Earth and Life Environmental Engineering Katsuya SATO · Associate Professor / Social Environment Systems Engineering, Ecosystem Engineering, Earth and Life Environmental Engineering

Shin-ichi Ito · Assistant Professor / Social Environment Systems Engineering, Ecosystem Engineering, Earth and Life Environmental Engineering

Target⟩ To understand the present environmental problems and to acquire the advanced of the technique and the approach to solve problems on the point of the ecosystem engineering.

Outline⟩ To explain the structure and function of environmental systems, the relationship between civil life and disaster prevention systems, and the changes of social system due to the technological revolution and the policy.

Style⟩ Lecture �������⟩ environmental policy, natural disaster, energy saving technology,

welfare technology, miromechanical engineering Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Requirement⟩ [要件] Notice⟩ [注意]

Goal⟩

1. To understand the factor of environmental systems

2. To understand the strucure of environmental systems

3. To understand the civil life and natural disaster prevention systems

Schedule⟩

1. The policy and natural disaster prevention measures (1)

2. The policy and natural disaster prevention measures (2)

3. The factor of environmental systems

4. Example of environmental systems 5. The structure of environmental systems

6. Example of structure of environmental systems

7. Civil life and nano-technology 1

8. Civil life and nano-technology 2

9. Midterm presentation

10. Well being technology 1

11. Well being technology 2

12. Ecosystem engineering 1

13. Ecosystem engineering 2

14. Hydrogen engine technology

15. Diesel engine engineering Evaluation Criteria⟩ Assignments count 100% Textbook⟩ To be introduced in th class Reference⟩ To be introduced in th class

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216598 Student⟩ [対象学生] Contact⟩

⇒ SATO (eco705, +81-88-656-2168, katsuyas@eco.tokushima-u.ac.jp) ����